# MITP-25-079

### v1

@@ 5 CKM matrix unitarity
it is necessary to have access to these three matrix elements. To this end, $|V_{ud}|$ can be extracted with high precision from super-allowed nuclear $\beta$-decays, while $|V_{ub}|$ can be neglected given its small value and the current precision for the other two matrix elements [ParticleDataGroup:2024cfk]. Lastly, we extract $|V_{us}|/|V_{ud}|$ through the use of Eq. 1.1, a lattice determination of $f_{K^{\pm}}/f_{\pi^{\pm}}$ – after including strong isospin-breaking effects in Eq. 4.16 – and taking into account the QED correction factor $\left(1+\delta_{\rm EM}^{K}-\delta_{\rm EM}^{\pi}\right)$. This allows us to extract $|V_{us}|$ by using the aforementioned value of $|V_{ud}|$.
@@
To compute $\delta_{\rm SU(2)}$, we use $f=113(28)$ MeV, $R=38.1(1.5)$ as reported by [FlavourLatticeAveragingGroupFLAG:2024oxs], resulting in
@@
where “stat” and $\chi-{\rm cont}$” correspond to the errors in Eq. 4.16, i.e. they are uncertainties related to the lattice data, and “SU(2)” corresponds to the uncertainty in Eq. 5.6. In Fig. 6 we compare our final estimate with results from other collaborations that enter the FLAG average for $N_{f}=2+1$ [FlavourLatticeAveragingGroupFLAG:2024oxs].
Figure 6: Comparison of our result for $f_{K^{\pm}}/f_{\pi^{\pm}}$ with strong isospin-breaking corrections, as given in Eq. 5.7, to other group’s results entering the FLAG average for $N_{f}=2+1$ [FlavourLatticeAveragingGroupFLAG:2024oxs]. CLQCD 23 refers to [CLQCD:2023sdb], QCDSF/UKQCD 16 to [QCDSF-UKQCD:2016rau], BMW 16 to [Durr:2016ulb], RBC/UKQCD 14B to [RBC:2014ntl], MILC 10 to [MILC:2010hzw], BMW 10 to [BMW:2010xmi] and HPQCD/UKQCD 07 to [Follana:2007uv]. The point ALPHA 25 corresponds to our previous determination of $f_{K}$ in [Bussone:2025wlf] based on an analysis setting the scale employing the intermediate scale $\sqrt{t_{0}}$.
For the QED correction $\delta_{\rm EM}^{\rm PS}$ to the experimental measurement of $\Gamma({\rm PS}\rightarrow l\overline{\nu}_{l}[\gamma])$, ${\rm PS}=\pi,K$, we reproduce the same steps as in [Marciano:1993sh, Cirigliano:2007ga], and we will argue for a conservative choice of the corresponding uncertainty. The $\delta_{\rm EM}^{\rm PS}$ factor reads
@@
where the first term is the universal long distance contribution to one-loop, assuming that the pseudoscalar state PS is point-like, while the terms proportional to $c_{i}^{\rm PS},\;\tilde{c}_{i}^{\rm PS}$ parameterise the structure-dependent part. To the current precision it is enough to consider the $c_{1}^{\rm PS}$ terms, [^6] Adding the missing terms, using the results in [Descotes-Genon:2005wrq, Cirigliano:2007ga], amounts to a shift of a $16\%$ of $\sigma$ in the central value w.r.t. Eq. 5.11. whose values are given, using large-$N_{c}$ methods, by [Ananthanarayan:2004qk, Descotes-Genon:2005wrq, Cirigliano:2011tm]
@@
where we took only the central values in Eq. 5.9 and then assumed 100% uncertainty for $\Delta c_{1}$, in order to cover the structureless pion and kaon limit in which $\Delta c_{1}=0$, which we consider to be the most conservative choice. Taking $l=\mu$ and with the function $F(x)$ explicitly given in [Marciano:1993sh], we get
@@
Using the experimental value [ParticleDataGroup:2024cfk]
@@
together with the QED correction in Eq. 5.11 and the values for $m_{K^{\pm}},\;m_{\pi^{\pm}},\;m_{\mu}$ reported in [ParticleDataGroup:2024cfk], following Eq. 1.1 we arrive at
@@
Employing the value of $|V_{ud}|$ from super-allowed $\beta$-decays $|V_{ud}|=0.97367(32)_{\beta{\rm-dec.}}$ [ParticleDataGroup:2024cfk] and neglecting $|V_{ub}|$, whose value squared is $\mbox{O}(10^{-5})$ [ParticleDataGroup:2024cfk] and thus negligible to the current precision, we obtain
@@
For completeness, we also test employing the $N_{f}=2+1+1$ lattice result $|V_{us}|=0.22328(58)_{|V_{us}|}$ determined from $f_{+}(0)$ [FlavourLatticeAveragingGroupFLAG:2024oxs] together with our result in Eq. 5.17 to determine $|V_{ud}|$, from which we obtain
@@
which is at $1.8\sigma$ from the Standard Model unitarity constraint.
If we repeat the same excersise but using the $N_{f}=2+1$ lattice result $|V_{us}|=0.22377(75)_{|V_{us}|}$ [FlavourLatticeAveragingGroupFLAG:2024oxs] we find
@@ 6 Conclusions
In this work we have presented results for $f_{K}/f_{\pi}$ in pure isoQCD, in addition to $|V_{us}|/|V_{ud}|$, where strong isospin-breaking and QED effects have been included, which allows to perform a test of the unitarity of the first row of the CKM matrix. Two important features of our construction are that (i) it is based on a combination of a Wilson unitary action and a mixed action with Wilson twisted mass valence fermions [Bussone:2025wlf, Bussone:2023kag], and (ii) that it is based on $f_{\pi}$ to set the scale. This latter point allows to avoid any systematic uncertainty related to the continuum-limit extrapolation of theory scales like $t_{0}$ or $w_{0}$. We quote as our main results
@@ Acknowledgments
The authors want to thank Alberto Ramos for the very useful discussions and critical reading of a previous version of the manuscript. We also want to thank Gregorio Herdoiza and Carlos Pena for help in the data generation, fruitful discussions and feedback on a previous version of the manuscript. We are grateful to our colleagues in the CLS initiative for sharing ensembles. We acknowledge PRACE for awarding us access to MareNostrum at Barcelona Supercomputing Center (BSC), Spain and to HAWK at GCS@HLRS, Germany. The authors thankfully acknowledge the computer resources at MareNostrum and the technical support provided by Barcelona Supercomputing Center (FI-2020-3-0026). We thank CESGA for granting access to Finis Terrae II. We also acknowledge support from the Spanish Research Agency (Agencia Estatal de Investigación) through national project CNS2022-13600, AEI/MCIU through grant PID2023-148162NB-C21 and ASFAE/2022/020.
Appendix A Appendix: QCD scheme variation and impact
In this Appendix we test the impact of modifying the definition of isoQCD by taking different values of $m_{\pi,K}^{\rm ph}$ and $f_{\pi}^{\rm ph}$ with respect to those in Eq. 1.2. In particular, we take the values given in [Aoki:2016frl]
@@ Appendix A Appendix: QCD scheme variation and impact
For completeness, we also quote the result for the derivatives of $f_{K}/f_{\pi}$ with respect to the input quantities in Table LABEL:tab:derivatives. These derivatives are estimated using the automatic differentiation tools implemented in the ADerrors package [Ramos:2018vgu].

### v2

@@ 5 CKM matrix unitarity
it is necessary to have access to these three matrix elements. To this end, $|V_{ud}|$ can be extracted with high precision from super-allowed nuclear $\beta$-decays, while $|V_{ub}|$ can be neglected given its small value and the current precision for the other two matrix elements [37]. Lastly, we extract $|V_{us}|/|V_{ud}|$ through the use of Eq. 1.1, a lattice determination of $f_{K^{\pm}}/f_{\pi^{\pm}}$ – after including strong isospin-breaking effects in Eq. 4.16 – and taking into account the QED correction factor $\left(1+\delta_{\rm EM}^{K}-\delta_{\rm EM}^{\pi}\right)$. This allows us to extract $|V_{us}|$ by using the aforementioned value of $|V_{ud}|$.
@@
To compute $\delta_{\rm SU(2)}$, we use $f=113(28)$ MeV, $R=38.1(1.5)$ as reported by [3], resulting in
@@
where “stat” and $\chi-{\rm cont}$” correspond to the errors in Eq. 4.16, i.e. they are uncertainties related to the lattice data, and “SU(2)” corresponds to the uncertainty in Eq. 5.6. In Fig. 6 we compare our final estimate with results from other collaborations that enter the FLAG average for $N_{f}=2+1$ [3].
Figure 6: Comparison of our result for $f_{K^{\pm}}/f_{\pi^{\pm}}$ with strong isospin-breaking corrections, as given in Eq. 5.7, to other group’s results entering the FLAG average for $N_{f}=2+1$ [3]. CLQCD 23 refers to [27], QCDSF/UKQCD 16 to [6], BMW 16 to [23], RBC/UKQCD 14B to [5], MILC 10 to [4], BMW 10 to [22] and HPQCD/UKQCD 07 to [24]. The point ALPHA 25 corresponds to the determination of $f_{K}$ in [12] based on an analysis setting the scale employing the intermediate scale $\sqrt{t_{0}}$.
For the QED correction $\delta_{\rm EM}^{\rm PS}$ to the experimental measurement of $\Gamma({\rm PS}\rightarrow l\overline{\nu}_{l}[\gamma])$, ${\rm PS}=\pi,K$, we reproduce the same steps as in [34, 15], and we will argue for a conservative choice of the corresponding uncertainty. The $\delta_{\rm EM}^{\rm PS}$ factor reads
@@
where the first term is the universal long distance contribution to one-loop, assuming that the pseudoscalar state PS is point-like, while the terms proportional to $c_{i}^{\rm PS},\;\tilde{c}_{i}^{\rm PS}$ parameterise the structure-dependent part. To the current precision it is enough to consider the $c_{1}^{\rm PS}$ terms, [^6] Adding the missing terms, using the results in [20, 15], amounts to a shift of a $16\%$ of $\sigma$ in the central value w.r.t. Eq. 5.11. whose values are given, using large-$N_{c}$ methods, by [1, 20, 14]
@@
where we took only the central values in Eq. 5.9 and then assumed 100% uncertainty for $\Delta c_{1}$, in order to cover the structureless pion and kaon limit in which $\Delta c_{1}=0$, which we consider to be the most conservative choice. Taking $l=\mu$ and with the function $F(x)$ explicitly given in [34], we get
@@
Using the experimental value [37]
@@
together with the QED correction in Eq. 5.11 and the values for $m_{K^{\pm}},\;m_{\pi^{\pm}},\;m_{\mu}$ reported in [37], following Eq. 1.1 we arrive at
@@
Employing the value of $|V_{ud}|$ from super-allowed $\beta$-decays $|V_{ud}|=0.97367(32)_{\beta{\rm-dec.}}$ [37] and neglecting $|V_{ub}|$, whose value squared is $\mbox{O}(10^{-5})$ [37] and thus negligible to the current precision, we obtain
@@
For completeness, we also test employing as input the value $|V_{us}|=0.22328(58)_{|V_{us}|}$, obtained from the experimental determination of $|V_{us}|f_{+}(0)$ [36] and the $N_{f}=2+1+1$ lattice result for $f_{+}(0)$ [3]. Using this input for $|V_{us}|$ together with our result in Eq. 5.17, we obtain
@@
which is at $1.8\sigma$ from the Standard Model unitarity constraint. If we repeat the same excersise but using now the $N_{f}=2+1$ lattice result for $f_{+}(0)$ [3], then $|V_{us}|=0.22377(75)_{|V_{us}|}$ and we find
@@ 6 Conclusions
In this work we have presented results for $f_{K}/f_{\pi}$ in pure isoQCD, in addition to $|V_{us}|/|V_{ud}|$, where strong isospin-breaking and QED effects have been included, which allows to perform a test of the unitarity of the first row of the CKM matrix. Two important features of our construction are that (i) it is based on a combination of a Wilson unitary action and a mixed action with Wilson twisted mass valence fermions [12, 13], and (ii) that it is based on $f_{\pi}$ to set the scale. This latter point allows to avoid any systematic uncertainty related to the continuum-limit extrapolation of theory scales like $t_{0}$ or $w_{0}$. We quote as our main results
@@ Acknowledgments
The authors want to thank Alberto Ramos for the very useful discussions and critical reading of a previous version of the manuscript. We also want to thank Gregorio Herdoiza and Carlos Pena for help in the data generation, fruitful discussions and feedback on a previous version of the manuscript. We are grateful to our colleagues in the CLS initiative for sharing ensembles. We acknowledge PRACE for awarding us access to MareNostrum at Barcelona Supercomputing Center (BSC), Spain and to HAWK at GCSHLRS, Germany. The authors thankfully acknowledge the computer resources at MareNostrum and the technical support provided by Barcelona Supercomputing Center (FI-2020-3-0026). We thank CESGA for granting access to Finis Terrae II. We also acknowledge support from the Spanish Research Agency (Agencia Estatal de Investigación) through national project CNS2022-136005, AEI/MCIU through grant PID2023-148162NB-C21 and ASFAE/2022/020.
Appendix A QCD scheme variation and impact
In this Appendix we test the impact of modifying the definition of isoQCD by taking different values of $m_{\pi,K}^{\rm ph}$ and $f_{\pi}^{\rm ph}$ with respect to those in Eq. 1.2. In particular, we take the values given in [2]
@@ Appendix A QCD scheme variation and impact
For completeness, we also quote the result for the derivatives of $f_{K}/f_{\pi}$ with respect to the input quantities in Table LABEL:tab:derivatives. These derivatives are estimated using the automatic differentiation tools implemented in the ADerrors package [40].